The number of solutions of the equation (1-2(log x2)2)/(log x-2 (logx )2) = 1

Grade:11

1 Answers

Chetan Mandayam Nayakar
12 years ago

Dear Kshitiz,

let logx=y, equation becomes (1-8y2)=y-2y2

6y2+y-1=0,y=(-1±5)/2=-3 or +2

log(x)=-3 or 2,x=e-3 or e2

Thus, there are two solutions

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Kshitiz !!!

 


Regards,

Askiitians Experts

Chetan Mandayam Nayakar